Your demand and supply functions are given by D: P=50-Q and S: P=10+3Q. Determine the market equilibrium price and quantity. If a price floor is put into place at P=30, determines the shortage or surplus, if any (think carefully). Draw a graph and show your calculations for full credit.
Answer #1

Answer : At market equilibrium, demand = supply.

Based on given information,

=> 50 - Q = 10 + 3Q

=> 50 - 10 = 3Q + Q

=> 40 = 4Q

=> Q = 40 / 4

=> Q = 10

Now from demand function we get,

P = 50 - 10

=> P = 40

Therefore, the equilibrium price is 40 and quantity is 10.  

Now if P = 30 then the market face a situation of shortage. Because the equilibrium price, P = 40 is higher than P = 30 and at lower price level the quantity demanded become higher than the quantity supplied. This is shown by the following picture's diagram. In the following picture's diagram the market faces excess demand or shortage at P = 30.